# Long Cheng1 Jiafei Duan1,211footnotemark: 1 Yi Ru Wang1 Haoquan Fang122footnotemark: 2 Boyang Li122footnotemark: 2 Yushan Huang1 Elvis Wang3 Ainaz Eftekhar1,2 Jason Lee1 Wentao Yuan1 Rose Hendrix2 Noah A. Smith1,2 Fei Xia1 Dieter Fox1 Ranjay Krishna1,2

### v1

@@ 4.4 What Other Factors Drive Pointing Performance?
Method Affordance Spatial Reasoning Steerability Counting Average
GPT-4o + In-Context (2-shots) 46.0 26.7 23.9 22.5 33.2 30.4
GPT-4o + Chain-of-Thought (CoT) wei2022chain 41.4 24.6 21.2 13.5 32.1 26.6
GPT-4o + Unparsed language instruction 37.8 23.7 19.7 21.9 31.1 26.9
GPT-4o (Default) 42.4 25.6 23.8 24.5 31.1 29.5
Gemini-2.5-Flash + Chain-of-Thought (CoT) wei2022chain 44.4 25.6 24.9 26.0 33.7 30.9
Gemini-2.5-Flash (Default) 58.6 47.7 43.2 35.0 49.7 46.8
@@ 5 Limitations, Discussion, and Conclusions
Limitations. The current annotation pipeline relies on the Segment Anything Model (SAM) kirillov2023segment to generate initial masks, which annotators refine through a grid-based interface. While efficient, this approach often results in coarse and imprecise boundaries, particularly for fine-grained or irregular shapes. This degrades segmentation quality and introduces noise into downstream evaluation. Additionally, as large multimodal models are frequently trained on publicly available datasets, static benchmarks such as Point-Bench are at increasing risk of becoming part of training data, reducing their effectiveness in evaluating generalization. Lastly, Point-Battle currently selects model pairs uniformly at random, which leads to uninformative comparisons—especially between models with large performance gaps—limiting the efficiency of the evaluation process.

### v2

@@ 5 Limitations, Discussion, and Conclusions
Limitations. The current annotation pipeline relies on the Segment Anything Model (SAM) Kirillov et al. (2023) to generate initial masks, which annotators refine through a grid-based interface. While efficient, this approach often results in coarse and imprecise boundaries, particularly for fine-grained or irregular shapes. This degrades segmentation quality and introduces noise into downstream evaluation. Additionally, as large multimodal models are frequently trained on publicly available datasets, static benchmarks such as Point-Bench are at increasing risk of becoming part of training data, reducing their effectiveness in evaluating generalization. Lastly, Point-Battle currently selects model pairs uniformly at random, which leads to uninformative comparisons—especially between models with large performance gaps—limiting the efficiency of the evaluation process.